OpenStudy (anonymous):

MEDAL AND FAN! :] Could someone help ME with 1 question?

OpenStudy (anonymous):

I will help you! so what is the question?

OpenStudy (anonymous):

Given the equation y - 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form. Answer Choices: -three fourthsx + y = 10 3x - 4y = -40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (the_fizicx99):

First solve it into the form of y = mx + b then convert it from y = mx + b into Ax + By = C Do you know how to solve?

OpenStudy (anonymous):

no i dont

OpenStudy (the_fizicx99):

kk no worries \(\ y - 4 = \frac{3}{4}(x+8) \) apply distributive property to \(\ \frac{3}{4}(x) + \frac{3}{4}(8) \)

OpenStudy (anonymous):

so the answer is 8?

OpenStudy (the_fizicx99):

No

OpenStudy (anonymous):

which do you think the best answer would be? :]

OpenStudy (the_fizicx99):

@MyChem what is \(\ \frac{3}{4}(x) \) ? I can't say which is the correct answer without solving the problem first, I'm solving it with you. :)

OpenStudy (anonymous):

idk :c

OpenStudy (anonymous):

@tHe_FiZiCx99 ya still here?

OpenStudy (the_fizicx99):

\(\ \frac{3}{4}(x) = \frac{3}{4}x \)

OpenStudy (anonymous):

?

OpenStudy (the_fizicx99):

Ok, do you know how to apply distributive property?

OpenStudy (anonymous):

ah no i just started this lesson...srry

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):